How many number of atoms does contain in 7g KOH(POTASSIAM HIDRO OXIDE)?

1
Expert's answer
2022-05-17T13:21:03-0400

M(KOH)=39+16+1=56 g/mole

n(KOH)=7/56=0.125 mole

1 mole contains 6.02x1023 molecules, then 0.125 mole contains 0.125x6.02x1023=0.7525 x1023 molecules. Rach molecule contains 3 atoms: K,O,H, so it is 0.7525x3x1023 =2.2575 x1023 atoms